1. m/12= 86
2.1/2m =6
3. 4z =16
4. 4-2/9 = 20/27
I dont under stand, I have alot however I need help to understand how.

Answer by stanbon(75887)

1.m/12= 86
m = 12*86
----------------
2.1/2m =6
m = 2*6
---------------
3. 4z =16
z = 16/4
--------------
4. 4-(2/9) = (20/27)
Multiply both sides by 27 to get:
4*27 - 2*3 = 20
28 - 6 = 20
22 = 20
I think something is missing in your post of this problem
=========================================================
Cheers,
Stan H.